MODEL STUDIES OF PARTIAL HYDRATION OF PROTONATED PEPTIDE IONS: I. THE FORMAMIDE-WATER CLUSTERS

\underline{C.-C. WU} and H.-C. CHANG, Institute of Atomic and Molecular Sciences, Academia Sinica, P. O. Box 23-166, Taipei, Taiwan 106, R.O.C..